The lines described by y=(a+12)x+3 and y=4ax are parrell. What is the value of a?
Answer by mananth(16946)

y=(a+12)x+3 and y=4ax are parallel. What is the value of a?
the slopes are same when the lines are parallel
(a+12)= 4a
3a=12
a=4